OpenStudy (becca9898):

Please help!! Kara is 4 more than twice as old as Julia. The sum of their ages is 13. How old is Julia?

imqwerty (imqwerty):

let julia's age be x kara;s age will be 2x+4 sum of thie ages will be -->x + (2x+4) its given that sum of ages is13 so x+2x+4=13 3x+4=13 3x=9 x=3 so julia is 3

OpenStudy (anonymous):

So simplify the word problem. that's \[x = \frac{ 13 - 4}{ 2}\] Find x.

OpenStudy (becca9898):

9/2 would be 4.5, that does not seem right. I think after going over my work the answer would have to be 3